OpenStudy (anonymous):

lim x->0 tan(2x)csc(5x)

OpenStudy (sirm3d):

\[\tan 2x\csc 5x=\frac{\sin 2x}{\cos 2x}\cdot\frac{1}{\sin 5x}=\frac{1}{\cos 2x}\cdot 2\frac{\sin 2x}{2x}\cdot \frac{1}{5}\frac{5x}{\sin 5x}\]

OpenStudy (raden):

alternative : lim x->0 tan(2x)csc(5x) = lim x->0 tan(2x)/sin(5x) use the basic formula : lim x->0 tan(Ax)/sin(Bx) = A/B